)A normally distributed population has mean 25.6 and standard deviation 3.3

(a) Find the probability that a single randomly selected element X of the 

population exceeds 30.

(b)Find the mean and standard deviation of X¯ for samples of size 9.

(C)Find the probability that the mean of a sample of size 99 drawn from this 

population exceeds 30.


1
Expert's answer
2021-01-08T11:48:37-0500

a) mean=25.6, sd=3.3

P(X>30)

Z="\frac{X-\mu} {\sigma}"

="\frac{30-25.6}{3.3}" =1.33

P(z>1.33)=0.09176 from the z tables.

=0.09176

b) mean of "\bar X"

"\mu=\bar X=25.6"

Standard deviation

"s=\frac{\sigma} {\sqrt n}"

="\frac {3.3}{\sqrt 9}" =1.1

c) P(X>30) given n=99

Z="\frac{30-25.6}{\frac{3.3}{\sqrt{99}}}" =13.27

P(z>13.27)"\approx" 0 from the standard normal tables.

=0